OpenStudy (lukecrayonz):

You and your colleagues have randomly chosen 15 children from the city schools to pilot your new reading program starting this January. An irate parent (in an editorial to the local newspaper) claims you are favoring males in your selection. Your random sample has 9 males and 6 females. (The entire population of children this age in your community, from which you have chosen your sample, is 70% male and 30% female.) 1.How likely is it to draw a sample of 15 with 6 or fewer females from this population?

OpenStudy (anonymous):

What is the prob of choosing all males?

OpenStudy (anonymous):

If we assume the entire population is really big

OpenStudy (lukecrayonz):

The parents seem to not like it @Redcan

OpenStudy (anonymous):

At least one editorial writing parent.

OpenStudy (anonymous):

Ok, hint #2... Are the events that you choose no females and exactly one female independent?

OpenStudy (lukecrayonz):

I would assume so

OpenStudy (anonymous):

I agree, you can't choose exactly one and exactly 0 at the same time

OpenStudy (anonymous):

So now we know, \[P(k\le 6)=P(k=0)+P(k=1)+\dots+P(k=?)\]